03-10-2009
All ISS systems continue to function nominally, except those noted previously or below. EVA-21A Day.

With today's wake period greatly lengthened from yesterday to ~20 hrs (3:30am – 11:30pm EDT), the crew continued on a sleep cycle shift protocol designed to adjust to the requirements of today's EVA-21A and the following Mission 15A docked period (see Wake/Sleep Schedule, below, based on a nominal STS-119 launch on Wednesday).

The Russian Orlan spacewalk by CDR Fincke & FE-1 Lonchakov went underway at 12:22pm EDT and is scheduled to run for an estimated 5h 45m (i.e., through ~6:07pm tonight).

CDR Fincke started the day with the daily download of the accumulated data of the SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment from his Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of another week-long session with SLEEP, his third. [To monitor the crewmember's sleep/wake patterns and light exposure, the crewmembers wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment's laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary “job jar” task list.]

After morning inspection, all pre-EVA activities proceeded smoothly and on schedule, starting out with Lonchakov & Fincke taking another MO-9 “Urolux” urine biochemistry test before breakfast. [A second session with the Urolux equipment will be conducted by both crewmembers tonight (~7:20pm/7:35pm) immediately after post-EVA station repress.]

Pre-EVA preparations by the crew included –

  • Closing external protective window shutters on SM (Service Module, #6,8,12,13,14) and JPM (JEM Pressurized Module),
  • Deactivating the amateur radio equipment in the FGB to prevent RF interference with the Orlans' wireless in-suit Tranzit-B radio telemetry system,
  • Transferring pre-selected food choices and food utensils to the FGB for FE-2 Magnus during her “lockout” in the FGB & USOS (US Segment),
  • Deactivating the DS-7A Smoke Detector #1 in the SM PkhO (Transfer Compartment),
  • Setting up the PSS Caution & Warning System in the FGB,
  • Deactivating the Vozdukh carbon dioxide (CO2) removal system,
  • Removing SM air ducts to enable RO-PkhO hatch closure,
  • Deactivating VN1 & VN2 air heaters,
  • Powering down ventilation fans in the SM (VK1, VK2, VK3, VK4, VKYu1, VKYu2),
  • Turning off the SRVK-2M condensate water processor, and
  • Powering off electrical food heaters in the SM galley.

In addition, CDR Fincke and FE-1 Lonchakov had about an hour reserved to complete setting up DC1 (Docking Compartment) and PkhO systems for pre-EVA mode.

Next steps by the crew were to –

  • Check out the Orlan-M spacesuits and their systems, as well as the suit interface control panels (BSS) in DC1 & PkhO,
  • Open the MPEV (Manual Pressure Equalization Valve) in the Node 1's overhead hatch to prevent negative pressure on the hatch after the EVA,
  • Retest the BK-3 primary & backup oxygen (O2) tanks of the Orlans and DC1,
  • Disassemble the DC1 air duct, but leaving the V3 fan in place, and
  • Set up the STTS communications/telemetry links necessary for the spacewalk from the DC1, including switching over the caution & warning system from the regular PSS console to the EVA support panel (POV). [Most activities were paced by RGS (Russian ground site) comm window passes.]

After a midday “snack” at ~8:00-8:40am, the CDR checked out the EVA cameras by “test firing” them and transferred them to the DC1 where the spacewalkers conducted final inspections of the suits, BSS interface units & biomedical parameter telemetry to RGS (~9:02am), including VHF/voice & biomedical electrode belt and *telemetry hookups via the BSS (later by the wireless in-suit Tranzit-B radio telemetry system) for vital signs and equipment monitoring.

FE-2 Magnus supported the preparations by assisting Mike & Yuri in suit donning plus Orlan, BSS, comm systems & biomedical parameter telemetry tests.

In the DC1 & PkhO, after the Orlan & BSS checkouts Lonchakov & Fincke donned the spacesuits and ancillary gear at ~10:00am, assisting each other, then closed the hatchways between SM RO/PkhO (Working Compartment/Transfer Compartment) and PkhO/SU (DC1 Transfer Vestibule) at ~10:40am, with Magnus supporting from inside the PkhO.

At ~11:20am, the spacewalkers sealed the Orlan backpacks, followed by Orlan & BSS controls checks. Final checkout of suits and their controls included testing for leak during successive stages of depressurization. [Pressure inside the Orlans was reduced to 0.42 at (6.2 psi). After suit purge, the spacewalkers had a 30-minute oxygen prebreathe period, as pressures between DC-1 and the PkhO were equalized and then further reduced.]

ISS attitude control authority, which was handed over to RS MCS (Motion Control System) at ~9:40am to keep attitude stable during the (reactive) DC1 airlock depressurization, was returned to US momentum management by CMGs (Control Moment Gyroscopes) at ~12:37pm. At ~1:00pm, Russian thrusters were disabled for the duration of the EVA until approximately 5:27pm, i.e., after the spacewalkers have left the SM large diameter area.

Sandy Magnus retreated into the FGB at ~11:30am and closed the hatch to the SM, which now serves as a backup airlock to the DC1. [The new US WHC (Waste & Hygiene Compartment) renders the use of the Soyuz ASU toilet facility unnecessary.]

The spacewalk began after a final leak check on the four BK-3 O2 tanks. At end of prebreathe, DC1 pressure was down to 15 mmHg (Torr), holding for 5 min for a final cabin leak check, followed by switching the Orlans to autonomous (battery) suit power (~12:18pm) and opening of EV hatch #1 at 12:22pm EDT.

As the spacewalk got underway, the FE-2 returned to the US Lab to collect the regular periodic US WRS (Water Recovery System) water sample from the RIP (Rack Interface Panel) and prepare it for return on 15A for chemical & microbial post-flight analysis on the ground. [WRS sampling & checkouts are being conducted for 90 days, i.e., every 4 days: WRS water hose (TOCA inflight analysis) & microbial bag sample (inflight bacterial visual enumeration plus archival for return on 15A), every 8 days: an archival water sample (return on 15A), and monthly: a TOCA bag sample from PWD (tested inflight).]

Afterwards, Sandra also performed the regular changeout of the TOCA (Total Organic Carbon Analyzer)'s waste water bag, then conducted the periodic WPA (Water Processor Assembly) sample analysis in the TOCA after first priming (filling) the TOCA water sample hose. Results were transferred to SSC-7 (Station Support Computer 7) via USB drive for downlink and the data were also logged for calldown. [The current procedure is a work-around for TOCA's failed catalyst.]

After concluding her 24-hour heart rate collection for the CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS) experiment at ~8:25am this morning, Sandy took off the HM2 (Holter Monitor 2) and Actiwatch for temporary stowage, then downloaded the collected data.

The FE-2 also performed the periodic deployment of four passive FMK (Formaldehyde Monitoring Kit) sampling assemblies in the Lab (at P3, below CEVIS) and SM (at the most forward handrail, on panel 307) for two days, to catch any atmospheric formaldehyde on a collector substrate for subsequent analysis on the ground. [Two monitors each are usually attached side by side, preferably in an orientation with their faces perpendicular to the direction of air flow.]

Later, tonight, Magnus will collect the periodic atmospheric samples in the center of the Lab, SM and “Kibo” JPM with the US GSCs (Grab Sample Containers), afterwards packing them for return on 15A. [The sampling will be sequenced with the third VOA (Volatile Organics Analyzer) run (4:15pm-5:30pm) for postflight data comparison.]

The FE-2 also had one hour set aside for more cargo prepacking for 15A.

After return and ingress from the EVA, with DC1 airlock repressurization from SM cabin air, the crew will open hatches and reenter the SM for their second MO-9 “Urolux” biochemical urine test.

Sandy Magnus will break out and set up the US PHS (Periodic Health Status) PHS equipment for the standard post-EVA exam, scheduled for Mike Fincke tomorrow morning immediately after wakeup. [The assessment used the AMP (Ambulatory Medical Pack), stethoscope, oral disposable thermometer and ABPC (Automatic Blood Pressure Cuff) from the ALSP (Advanced Life Support Pack). All data were then logged on the MEC (Medical Equipment Computer) and the hardware stowed. On the MEC laptop, the PHS exam is guided by special IFEP (In-Flight Examination Program) software.]

With SM pressure restored back to normal, Magnus will then re-close the MPEV in the Node-1 overhead hatch and assist the spacewalkers in their post-EVA closeout activities.

After a joint “midnight” snack (~8:00pm), the crew will have another two hours for reconfiguring the ISS interior for normal operation, including resetting STTS communications, activating turned-off systems, and restoring systems configurations in the DC1 and SM to pre-EVA conditions, then installing the DC1 air ducts. Also, Mike & Yuri are to complete Orlan water feed line drying and other suit closeout ops.

Following a late dinner and food preparations for tomorrow, bedtime for the crew will commence at ~11:30pm EDT.

Sandy Magnus was scheduled for her regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer and ARED (Advanced Resistive Exercise Device).

At ~4:10am EDT, the crew joined in a Russian PAO TV exchange with Roskosmos management and selected high school students in the Open Lesson “Youth Chooses Space”. [The Lesson was conducted as part of the Federal Space Agency's project dedicated to 75th birthday anniversary (3/9) of Yuri Alekseyevich Gagarin, the first cosmonaut on the planet. Attendees included the heads of Roskosmos, StNIIMash and TsUP, Gagarin era cosmonauts Volynov Boris Valentinovich & Popov Pavel Romanovich, and the high school student winners of the 16th International Space Olympiad. The lesson was anchored by Anatoly Nikolaevich Perminov, the head of Roskosmos, and Irina Rossius, the host of Vesty TV channel.]

ISS Crew Sleep Shift Planning: To synchronize the ISS crew's timeline with today's EVA and STS-119/15A arrival & docked period, the station wake/sleep cycle is undergoing a number of shifts which started on 3/7. For the next few days, the schedule is as follows:


3/11 Wake: 9:00am – 12:30am 3/12
3/12 Wake: 9:00am – 2:00am 3/13
3/13 Wake: 10:30am – 1:50am 3/14
3/14 Wake: 10:20am – 1:20am 3/15
3/15 Wake: 9:50am – 12:50am 3/16
3/16 Wake: 9:20am – 12:50am 3/17
3/17 Wake: 9:20am – 12:20am 3/18
3/18 Wake: 8:50am – 11:50pm
3/19 Wake: 8:20am – 11:20pm
3/20 Wake: 7:50am – 11:20pm
3/21 Wake: 7:50am – 10:50pm
3/22 Wake: 7:20am – 10:50pm
3/23 Wake: 7:20am – 10:10pm
3/24 Wake: 6:40am – 9:30pm


EVA-21A Timeline: Today's Orlan EVA-21A by Lonchakov (EV1) & Fincke (EV2) began at ~12:22pm EDT (DC1 EV hatch open), to last an estimated 5h 45m, i.e., concluding at around 6:07pm. Objectives of the EVA (all fallen off the previous EVA-21 timeline) are –

  • Mount the EXPOSE-R hardware on the URM-D (Portable Multipurpose Workstation) on the SM RO l.d., connect it to the PF-3 connector patch panel and remove protective cover;
  • Photograph the URM-D with EXPOSE-R monoblock & cables, ROBOTIC hardware, IPI-SM hardware and routed cables;

  • Remove fasteners (Aramide straps) in the installation areas of the docking target and AR-VKA & 2AR-VKA antennas on DC1;

  • Close MLI (Multi-Layer Insulation) flap on the SM PF-10 connector patch panel;

  • Re-install SKK #9 removable cassette container in nominal position on SM;

  • Inspect & photograph Progress antenna ASF1-M-VKA from DC handrail 3034; and

  • Inspect & photograph conditions of ISS RS exterior & structural elements (“Panorama-2009” DTO).


STS-119/Discovery: Countdown for Discovery's launch on Wednesday (9:20pm EDT) is proceeding smoothly. Weather is forecast to be favorable for launch with only a slight chance (~10%) of a launch-prohibiting low-cloud ceiling.

15A Crew & Mission Timeline:

  • CDR: Lee Archambault
  • PLT: Dominic Antonelli

  • MSs: Joseph Acaba; John Phillips; Steven Swanson; Richard Arnold

  • ISS FE-2: Koichi Wakata (UP); Sandra Magnus (DOWN)

  • FD1 (3/11) -- Launch 9:20pm EDT

  • FD2 (3/12) -- TPS inspection using OBSS; checkout EVA suits; prepare for rendezvous/docking

  • FD3 (3/13) -- Rendezvous; RPM, docking (~6:27pm); exchange Soyuz seat liners

  • FD4 (3/14) -- Unberth S6 truss w/SSRMS; handoff S6 to SRMS; move SSRMS to WS1, hand S6 back to SSRMS; park SSRMS with S6 at overnight position; prepare for EVA1; campout (Swanson & Arnold)

  • FD5 (3/15) -- EVA1; install S6 truss & solar arrays (~3 pm–9:30pm)

  • FD6 (3/16) -- Focused TPS inspection with OBSS on SRMS (if not required, deploy solar array wings); prepare for EVA2; campout (Swanson & Acaba)

  • FD7 (3/17) -- EVA2; prepare P6 battery R&R (Mission 2JA); P1/P3 tasks; deploy P3 UCCAS & S3 PAS (~2:15pm–8:45pm)

  • FD8 (3/18) -- Deploy two S6 solar array wings (115 ft long); move MT from WS4 to WS1; prepare for EVA3; campout (Arnold & Acaba)

  • FD9 (3/19) -- EVA3; relocate CETA; lube SPDM LEE B; replace two RPCMs; S1 tasks (~1:15pm–7:45pm)

  • FD10 (3/20) -- Crew off duty; continue cargo transfers; joint news conference; prepare for EVA4; campout (Swanson & Arnold)

  • FD11 (3/21) -- EVA4; JEM GPS antenna; reconfigure Z1 patch panel; photograph thermal radiators; S3 PAS; WETA (~12:45pm–7:15pm)

  • FD12 (3/22) -- Crew off duty; final cargo transfers; reboost; close hatches (~3:27pm)

  • FD13 (3/23) -- Undock (~10:23am); flyaround; late TPS inspection using OBSS

  • FD14 (3/24) -- Orbiter FCS checkout, RCS hot fire

  • FD15 (3/25) -- Nominal deorbit (2:24pm); landing (3:27pm KSC).



No CEO photo targets uplinked for today.
CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov <http://eol.jsc.nasa.gov/> (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 8:02am EST [= epoch])
Mean altitude -- 354.9 km
Apogee height -- 361.4 km
Perigee height -- 348.5 km
Period -- 91.64 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0009624
Solar Beta Angle -- 45.8 deg (magnitude increasing)
Orbits per 24-hr. day -- 15.71
Mean altitude loss in the last 24 hours -- 42 m
Revolutions since FGB/Zarya launch (Nov. 98) -- 59042

Significant Events Ahead (all dates Eastern Time, some changes possible!):
03/11/09 -- STS-119/Discovery/15A launch – S6 truss segment --- 9:20:10pm EDT
03/13/09 -- STS-119/Discovery/15A docking --- 6:27pm EDT
03/23/09 -- STS-119/Discovery/15A undocking --- 10:23am EDT
03/25/09 -- STS-119/Discovery/15A deorbit (Orbit 217, 2:24pm) & landing --- 3:27pm EDT (KSC)
03/26/09 -- Soyuz TMA-14/18S launch (7:49am EDT)
03/28/09 -- Soyuz TMA-14/18S docking (DC1; 9:15am EDT)
04/07/09 -- Soyuz TMA-13/17S undocking & landing
05/06/09 -- Progress 32P undocking & deorbit (under review)
05/07/09 -- Progress 33P launch
05/12/09 -- STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
05/12/09 -- Progress 33P docking
05/15/09 -- STS-127/Endeavour/2J/A launch - JEM EF, ELM-ES, ICC-VLD
05/27/09 -- Soyuz TMA-15/19S launch
05/29/09 -- Soyuz TMA-15/19S docking (FGB nadir)
Six-person crew on ISS
07/17/09 -- Progress 33P undock & deorbit
07/20/09 -- Soyuz TMA-14/18S relocation (to DC1)
07/24/09 -- Progress 34P launch
07/26/09 -- Progress 34P docking (SM aft)
08/06/09 -- STS-128/Discovery/17A – MPLM (P), LMC
09/01/09 -- H-IIB (JAXA HTV-1) -- tentative
11/10/09 -- Soyuz 5R/MRM2 (Russian Mini Research Module, MIM2) on Soyuz -- tentative
11/12/09 -- STS-129/Atlantis/ULF3 - ELC1, ELC2
12/10/09 -- STS-130/Endeavour/20A – Node-3 + Cupola -- tentative
02/11/10 -- STS-131/Atlantis/19A – MPLM(P), LMC -- tentative
04/08/10 -- STS-132/Discovery/ULF4 – ICC-VLD, MRM1 -- tentative
05/31/10 -- STS-133/Endeavour/ULF5 – ELC3, ELC4 -- tentative
12/XX/11 -- Proton 3R/MLM w/ERA.